OpenStudy (anonymous):

Does anyone have any advice for converting a number in tangent? Example, for arctan(1), its pi/4. What is it for arctan(1/2) ?or any other number?

hartnn (hartnn):

for what angle x is tan x = 1/2 that angle is the answer to arctan(1/2)

hartnn (hartnn):

if its a non standard angle, i just use a calculator or google lol

OpenStudy (anonymous):

yes , but I want to indtroduce pi in the number:p like for arctan(1).

hartnn (hartnn):

arctan (1/sqrt 3) = 30 degrees, to convert degrees in to radians, you multiply by \(\pi/180\) so, 30 degrees = \(30\pi/180 =\pi/6 \)

OpenStudy (anonymous):

sqrt?

hartnn (hartnn):

\(sqrt = \sqrt {}\)

OpenStudy (anonymous):

and how did you know that it was 30 degrees?

hartnn (hartnn):

because 30 is a standard angle and i knew tan 30 = 1/ sqrt 3

hartnn (hartnn):

or i can deduce it tan 30 = sin 30/cos 30 = 1/2 / sqrt 3/2 = 1/ sqrt 3

OpenStudy (anonymous):

Okay thank you :)

hartnn (hartnn):

for all non-standard angles, use calculator :) if your calculator is in radian mode, you will get answer in terms of pi, else if t is in degree mode, you will get answer in degrees and you need to convert it to radians as i did for 30 degrees

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

but its like: If My calculator is in radian, when I do : arctan(1/2) , then I /180 and it givess me the answer with pi ?

hartnn (hartnn):

arctan(1/2) = 0.4634 radians \(= (0.4634/\pi) \pi ~ radians \)

hartnn (hartnn):

= \(0.1475\pi^r\)

OpenStudy (anonymous):

ohhhh okkkkk Now I understaand for real hihi :)

hartnn (hartnn):

i am happy that you understood :)